є0nsider the following abbreviated financial statements for Parrothead Enterprises: PARROTHEAD ENTERPRISES PARROTHEAD ENTERPRISES 2014 and 2015 Partial Balance Sheets 2015 Income Statement Liabilities & OwnersSal Assets 12728 Equity 2015 2014 2015 2014 Costs 5005 Current assets 968 961 Current liabilities 333 495 Depreciation 1065 Net fixed assets 3523 4044 Long-term debt 2052 2187 Interest 190 What is owners equity?
Answer #1

SEE THE SCREENSHOT. ANY DOUBTS, HAPPY TO HELP YOU. THANK YOU. THUMBS UP PLEASE.

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard BD190 BA Font Alignment Number Styles Cells Edting BC BD BE BF BG BH BI BJ 173 174 175 176 PARROTHEAD 2015 2014 2015 2014 CURRENT ASSETS 968 961 CURRENT LIABILITIES LONG TERM DEBT 495 2187 23852682 178 179 180 181 182 183 184 185 186 187 188 189 190 191 2052 NET FIXED ASSETS 3523 4044 TOTAL LIABILITIES 2106 4491 OWNERS EQUITY 2323 TOTAL ASSETS 4491 5005 TOTAL LIABILITIES& 5005 OWNER S EQUITY OWNERS EQUITY = TOTAL ASSETS-TOTAL LIABILITIES FOR 2015 4491 - 2385 FOR 2014 = 5005-2682 = 2106 2323 , CASH BUDGET .. bud BV MV ratio VARIANCE BEP, OL FL ratios B-S t loss SALES BUDGET DİFF ANALYSIS overhead float | ORDER E14 erences: EV23 05:22 19-01-2019